# Lanczos-Pascal approach to correlation functions in chaotic quantum systems

### v2

@@ Appendix A The Lanczos-Pascal scheme as a derivative in Krylov space
The applicability of the Lanczos-Pascal scheme is closely related to the structure of the wave vector $\vec{\boldsymbol{\varphi}}$. Assuming that $\vec{\boldsymbol{\psi}}$ captures the first $N$ of $\vec{\boldsymbol{\varphi}}$ precisely, i.e. $\varphi_{n}(t)=\psi_{n}(t)$ for $n=0,...,N-1$, combining Eq. (3) and Eq. (6) we may write
00 0 $=\left(\textbf{S}-\textbf{L}\right)\vec{\boldsymbol{\varphi}}\ \Longleftrightarrow\ \vec{s}\cdot\vec{\boldsymbol{\varphi}}=0.$ (16)
Consequently, the existence of a matrix S that generates the first $N$ entries of $\vec{\boldsymbol{\varphi}}$ correctly is equivalent to a vector $\vec{s}$, orthogonal to the wave function. Further, with the vector $\vec{s}$ as defined in Eq. (5) the scalar product represents a discrete spatial derivative of the wave function. Hence whenever the wave vector $\vec{\boldsymbol{\varphi}}$ with suitably smooth with respect to its components we expect the LP scheme to yield appropriate results. In this interpretation the quantity $D$ corresponds to the order of the derivative. See Fig. 4 for an illustration of the wave vector $\vec{\boldsymbol{\varphi}}$ related to the fast-mode observable in the tilted-field Ising chain with $B_{z}=0.5$ for an impression of a smooth wave vector.
@@
Figure 4: Wave vector $\vec{\boldsymbol{\varphi}}$ related to the fast-mode of energy-density wave in the tilted-field Ising model with $B_{z}=0.5$ for various instances of time. Differently to the analysis in the main text, for simplicity we resort to solving the matrix equation (3).
While this consideration yields an intuitive picture of the inner workings of the LP method, it clearly does not provide solid justification for the applicability of the latter. This is so because the smoothness of $\vec{\boldsymbol{\varphi}}$ is phenomenologically demonstrated in Fig. 4, but not generally mathematically established based on properties of $b_{n}$. The latter will be done below.
Appendix B Smoothness of the wave vector
Consider the Laplace transform of Eq. (3) to find
@@ Appendix B Smoothness of the wave vector
where by $\widetilde{\varphi}_{k}(0)$, subsequently abbreviated by $\widetilde{\varphi}_{k}$, we denote the Laplace transform of the $k$th entry of the approximated wave vector $\vec{\boldsymbol{\varphi}}$ evaluated at zero. Note that $\widetilde{\varphi}_{0}$ equals the area under the curve, i.e, the quantity that we aim to estimate by $\mathcal{A}_{r}$ in Eq. (4). Equation (17) may be solved ”from top to bottom”. Doing so, every $\widetilde{\varphi}_{n}$ with $n$ even is determined up to a multiplicative factor $\widetilde{\varphi}_{0}$. Concretely,
@@
$\widetilde{\varphi}_{2n+1}$ $=\frac{1}{b_{1}},\frac{b_{2}}{b_{1}b_{3}},\frac{b_{2}b_{4}}{b_{1}b_{3}b_{5}},\dots,$ (19)
Eventually, we may write
@@
Consequently, convergence of $\mathcal{A}_{r}$ implies $\mathcal{A}_{r}\approx\widetilde{\varphi}_{0}$. Thus, for $r$ in the regime where $\mathcal{A}_{r}$ has converged, say at $r\geq R$, the square root term in Eq. (20) has to be very close to 1, independent of the specific $r$. This in turn implies that the sequence of the components $\widetilde{\varphi}_{r}$ has become smooth. This eventually amounts to the Laplace transformation of Eq. (16) at zero, i.e.
$\vec{s}\cdot\vec{\widetilde{\boldsymbol{\varphi}}}=0,$ (21)
typically being approximately fulfilled for pertinent $\vec{s}$, as this construction amounts to taking a discrete derivative of order $D$ of a smooth sequence. Further, we extend our arguments to nonzero arguments of the Laplace transformed quantities. Since the $\tilde{\varphi}_{n}(s)$ do not feature any poles on the right half complex plain, we may restrict our analysis to the imaginary axis, i.e. $\omega\neq 0$. (The Laplace transform thus amounts to a Fourier transform). On the latter, the Laplace/Fourier transform of (3) becomes
$-\vec{\boldsymbol{\varphi}}(0)=\left(-i\omega\ \mathbb{1}+\textbf{L}\right)\vec{\widetilde{\boldsymbol{\varphi}}}(\omega)$ (22)
@@
Upon defining
@@
we may link Eq. (23) to the setting where $\omega=0$ and write
@@
$\Leftrightarrow\left(f_{n+1}-f_{n-1}\right)$ $=i\omega f_{n}\left(\frac{\widetilde{\varphi}_{n+1}}{\widetilde{\varphi}_{n}}b_{n+1}\right)^{-1}.$ (26)
As argued around Eq. (20), for $n\geq R$, we have $\widetilde{\varphi}_{n+1}/\widetilde{\varphi}_{n}\approx 1$. As argued around Eq. (15), the $b_{n}$ form a smooth sequence in the same regime. With this, we take the continuum limit of Eq. (26) to find
@@
$\Rightarrow f(n)$ $=\exp\left(\frac{i}{2}\omega\int_{n_{0}}^{n^{\prime}}\frac{1}{b(n^{\prime})}\text{d}n^{\prime}\right).$ (28)
Figure 5: LP method in the tilted field Ising model (TFIM). The two-local observable $\mathcal{B}$ (left) and energy current $\mathcal{J}_{E}$ (right) in the tilted-field Ising model, given by Eq. (29) and (30) respectively. From top to bottom: Autocorrelation function obtained by the LP method and DQT (inset: log-plot), the correlation between the first two Krylov states of the respective observable, again by means of the LP method and DQT (inset: log-plot), the area estimate $\mathcal{A}_{r}$ with convergence position $R$ and the respective Lanczos coefficients $b_{n}$.
Considering finite frequencies $\omega\neq 0$, we infer from Eq. (28) that for Lanczos coefficients $b_{n}$ that grow according to the OGH, for small frequencies $\omega$, the functions $f_{n}$ oscillate slowly. Hence this allows to connect the picture with the $\omega=0$ setting and therefore implies that the LP scheme is also applicable in this regime. Conversely, for Lanczos coefficients that do not grow, the area $\mathcal{A}_{r}$ is constant for all $r$. However, it is evident from Eq. (28) that already at small frequencies the functions $f_{n}$ will oscillate strongly and therefore the scenario may not simply be mapped to the $\omega=0$ case. This behavior is perfectly in line with the LP technique explained in the main part of the paper. In the case of constant Lanczos coefficients the autocorrelation function is known to be a Bessel function, see [37], and further there is no such $D$ fulfilling $(b_{1}/b_{R+1})^{D}<1$ for any $R$.
Appendix C Further examples in the tilted-field Ising model
@@
$\mathcal{B}$ $\propto\sum_{k}J\sigma_{k}^{x}\sigma_{k+1}^{x}-\sigma_{k}^{z},$ (29)
$\mathcal{J}_{E}$ $=B_{x}\sum_{k}\sigma_{k}^{y}\left(\sigma_{k+1}^{z}-\sigma_{k-1}^{z}\right),$ (30)
with $J=1.05$, for which we compute $n_{\max}=40$ and $n_{\max}=44$ Lanczos coefficients respectively. In Fig. 5 we depict the LP curves for both observables for both $\psi_{0}(t)$ and $\psi_{1}(t)$ as well as the area estimator $\mathcal{A}_{R}$. For the two-local observable $\mathcal{B}$ we find a suitably converged area at $R=18$ and a necessary order $D=2$. For the energy current we have $(R,D)=(35,1)$. Again, the LP scheme gives rise to approximation that are distinguishable from the DQT data to the naked eye. Here, the functions $\psi_{0}(t)$ and $\psi_{1}(t)$ both only consist of at most 20 and 36 exponentially damped constituents for the two observables respectively.

### v3

@@ Appendix A Applicability of the Lanczos-Pascal scheme under the assumption of smooth $\varphi_{n}(t)$
The applicability of the Lanczos-Pascal scheme is closely related to the structure of the wave vector $\vec{\boldsymbol{\varphi}}$, see Fig. 4 for a sketch of a generic example.
In order to motivate the effective truncation of the Lanczos method, imposed by the Lanczos-Pascal scheme we start by assuming that there exists a set of (so far undetermined) coefficients $s_{n}$ such that
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{n=0}^{N}s_{n}\varphi_{n}(t)=0}$ (15)
holds true at all times $t$. In the following we drop the arguments for better clarity. Without loss of generality, we may set $s_{N}=1$. Rewriting yields
$\sum_{n=0}^{N-1}s_{n}\varphi_{n}+\varphi_{N}=0\Longleftrightarrow\varphi_{N}=-\sum_{n=0}^{N-1}s_{n}\varphi_{n}.$ (16)
This amounts to a guess on $\varphi_{N}$ based on all $\varphi_{n}$ with n<N1n<N-1 . The latter still fulfill the discrete Schrödinger equation (2), whereas for $\varphi_{N-1}$ we have
$\partial_{t}\varphi_{N-1}$ $=b_{N-1}\varphi_{N-2}-b_{N}\varphi_{N}$ (17)
$=b_{N-1}\varphi_{N-2}+b_{N}\sum_{n=0}^{N-1}s_{n}\varphi_{n}.$ (18)
Hence the Eqs. (2) and (18) form a linear, autonomous set of differential equations for $\varphi_{0},\dots,\varphi_{N-1}$. These equations may be cast into matrix form given in Eq. (6). Thus, if Eq. (15) would strictly apply, the LP scheme would produce the exact correlation function. As in practice we typically never strictly have Eq. (15), we employ the notation $\vec{\boldsymbol{\psi}}$ for the wave vector generated by S in the main text.
@@
Figure 4: Wave vector $\vec{\boldsymbol{\varphi}}$ related to the fast-mode of energy-density wave in the tilted-field Ising model with $B_{z}=0.5$ for various instances of time. Differently to the analysis in the main text, for simplicity we resort to solving the matrix equation (3). The curves are interpolated as the $\varphi_{n}(t)$ are only defined for $n\in\mathbb{N}$.
Eventually, the LP scheme imposes an explicit form for the coefficients $s_{n}$, where $n=0,\dots,N-1$. The value for $N$ is set by the parameters $R,D$ as $N=R+D$. With these $s_{n}$ we revisit the left-hand side of Eq. (16). With the choice of the $s_{n}$ from Eq. (5) we effectively implement a numerical ”derivative of the wave vector $\vec{\boldsymbol{\varphi}}$ with respect to nn” via a (possibly higher order) Newtonian difference quotient. The parameters $R,D$ determine both position and order of this derivative as
$\sum_{n=0}^{N-1}s_{n}\varphi_{n}+\varphi_{N}=\sum_{n=R}^{N-1}(-1)^{N+n}\binom{D}{n-R}\varphi_{n}+\varphi_{N},$ (19)
implementing a spatial derivative of order $D$ at position $R+D/2$. Consulting a typical example in Fig. 4 suggests that such an approach may be promising and plausible for sufficiently large $R,D$.
For the sake for clarity we will consider the setting addressed in Fig. 4 where of $R=20$, $D=2$ such that
$\{s_{n}\}_{n=0}^{N-1}=\{\underbrace{0,\dots,0}_{20},1,-2\},$ (20)
and eventually
$\sum_{n=0}^{N-1}s_{n}\varphi_{n}+\varphi_{N}=0=\varphi_{20}-2\varphi_{21}+\varphi_{22}.$ (21)
Consequently, the construction of the LP scheme exploits the spatial smoothness of the respective wave vector $\vec{\boldsymbol{\varphi}}$, effectively realizing a numerical derivative via a Newtonian difference quotient.
Appendix B Conditions on the $b_{n}$ for the smoothenss of the $\varphi_{n}(t)$
In this section, we examine to what extent smoothness of the wave vector can be expected, i.e. to what extend the validity of Eq. (15) with the usage of the respective Pascal numbers may be expected. To this end, we consider the Laplace transformation of $\vec{\boldsymbol{\varphi}}(t)$ which we denote by $\vec{\widetilde{\boldsymbol{\varphi}}}(s)$, where $s=\sigma+i\omega$. It is straightforward but important to realize that the validity of Eq. (15) at all times $t$ is equivalent to the validity of its Laplace transformed counterpart $\sum_{n=0}^{N}s_{n}\widetilde{\varphi}_{n}(s)=0$ at all $s$. Hence the LP scheme is expected to work if all $\vec{\widetilde{\boldsymbol{\varphi}}}(s)$ are smooth at sufficiently large $n$ in the sense of the previous section. First, we turn to the setting where $s=0$, where we can write the Laplace transformation of Eq. (3) as
@@ Appendix B Conditions on the $b_{n}$ for the smoothenss of the $\varphi_{n}(t)$
having used that $\varphi_{n}(0)=\delta_{n0}$. In the following we abbreviate $\widetilde{\varphi}_{n}(0)$ by $\widetilde{\varphi}_{n}$. The form of Eq. (22) allows to extract $\widetilde{\varphi}_{n}$ in terms of the Lanczos coefficients, solving the equation ”from top to bottom”. Concretely, we may rewrite Eq. (22) as
$-\delta_{n0}=b_{n}\widetilde{\varphi}_{n-1}-b_{n+1}\widetilde{\varphi}_{n+1},$ (23)
where $b_{0}=0.$ Up to a multiplicative constant $\widetilde{\varphi}_{0}$ we can solve iteratively to find the eventual entries as
@@
$\widetilde{\varphi}_{2n+1}$ $=\frac{1}{b_{1}},\frac{b_{2}}{b_{1}b_{3}},\frac{b_{2}b_{4}}{b_{1}b_{3}b_{5}},\dots.$ (25)
The quantity $\widetilde{\varphi}_{0}=\int_{0}^{\infty}\text{d}t\ \varphi_{0}(t)$ constitutes the area under the curve of the autocorrelation function and is in itself a well-studied quantity, see e.g. [33, 32, 34, 44]. Resorting to the estimator $\mathcal{A}_{r}$ of that quantity first mentioned in Ref. [44] and utilized in the main text (Eq. (4))we may work out a rewriting of the former as
@@
As now after some $r\geq R$ the quantity $\mathcal{A}_{r}$ converges, i.e. $\mathcal{A}_{r}\approx\widetilde{\varphi}_{0}$, we infer that in turn the vector $\vec{\widetilde{\boldsymbol{\varphi}}}$ needs to become smooth, so that thereafter, independently of $r$, the expression in the square root in Eq. (26) needs to become suitably close to 1. Note that while the convergence of the estimator $\mathcal{A}_{r}$ itself has been set into relation with the growth of Lanczos coefficients already in Ref. [44], the argument above directly draws a connection to structural properties of the wave vector. We now turn to non-zero arguments $s=\sigma+i\omega$ of the Laplace transform $\vec{\widetilde{\boldsymbol{\varphi}}}(s)$. Since the Laplace transform exists for Re(ss)>0>0, the $\widetilde{\varphi}_{n}(s)$ do not have poles in the right half of the complex plane. Hence we may restrict our analysis to the imaginary axis, $s=i\omega$, by the identity theorem. In this setting, the Laplace transform amounts to the Fourier transform. Concretely, for Eq. (3) we may write
$-\vec{\boldsymbol{\varphi}}(0)=\left(-i\omega\mathbb{1}-\textbf{L}\right)\vec{\widetilde{\boldsymbol{\varphi}}}(\omega),$ (27)
@@
We now aim to relate this setting to the $\omega=0$ scenario. Therefore we define
@@
such that inserting into Eq. (28) yields
@@
$\Leftrightarrow\left(f_{n+1}-f_{n-1}\right)$ $=i\omega f_{n}\left(\frac{\widetilde{\varphi}_{n+1}}{\widetilde{\varphi}_{n}}b_{n+1}\right)^{-1},$ (31)
having used that the functions $\widetilde{\varphi}_{n}(0)$ fulfill
$0$ $=b_{n+1}\widetilde{\varphi}_{n+1}-b_{n}\widetilde{\varphi}_{n-1}$ (32)
$\Leftrightarrow\widetilde{\varphi}_{n-1}$ $=\frac{b_{n+1}}{b_{n}}\widetilde{\varphi}_{n+1}.$ (33)
Resorting to the arguments above that suggest a smoothening of the wave vector at $s=0$, we argue that for some $n\geq R$, we have $\widetilde{\varphi}_{n+1}/\widetilde{\varphi}_{n}\approx 1$. In the continuum limit, i.e. $f_{n}\rightarrow f(n)$ and $b_{n}\rightarrow b(n)$, Eq. (31) may hence be written as
@@
which allows for the solution
$f(n)=\text{exp}\left(\frac{i}{2}\omega\int_{n_{0}}^{n}\frac{1}{b(n^{\prime})}\text{d}n^{\prime}\right)$ (35)
to the function relating the settings where $\omega=0$ and $\omega\neq 0$, for $n>n_{0}$ and $n_{0}\geq R+D$.
Here, the influence of the growth of the Lanczos coefficients becomes visible. Given that the Lanczos coefficients $b_{n}$ grow slowly or even are constant, already small frequencies $\omega$ will result in highly oscillatory behavior of the $f(n)$, compromising the mapping between the $\omega=0$ and $\neq 0$ settings. Conversely, for regularly, e.g. linearly growing Lanczos coefficients, the arguments presented for $\omega=0$ can similarly be made for small $\omega$, including oscillatory behavior into our analysis. These lines of reasoning can most prominently be seen at the example of the toy model, see Fig. 1. There, the slowness of the growth of the Lanczos coefficients ($\alpha=0.1$) evidently necessitates a higher value for $D$, as for $D=1$ the respective LP curve clearly fails to capture the actual dynamics, whereas in the quickly growing scenario ($\alpha=1$) already the order $D=1$ suffices to approximate the dynamics.
SUPPLEMENTAL MATERIAL
Details on dynamical quantum typicality
In this section, we provide more detail on the numerical method based on the dynamical quantum typicality (DQT), which we use to calculate the autocorrelation functions shown in the main text.
Let us start from a normalized Haar-random state in the Hilbert space
$|\psi\rangle=\sum_{j}C_{j}|E_{j}\rangle,$ (S1)
where $C_{j}$ are complex numbers whose real and imaginary parts are drawn independently from a Gaussian distribution. According to DQT, the autocorrelation function at infinite temperature $\beta=0$, $C(t)=\frac{1}{D}\text{Tr}[{\cal O}(t){\cal O}]$ can be written as
$C(t)=\langle\psi|e^{iHt}{\cal O}e^{-iHt}{\cal O}|\psi\rangle+\varepsilon(t).$ (S2)
The error scales as
$\varepsilon(t)\sim\frac{1}{\sqrt{D}}$ (S3)
with $D$ being the Hilbert space dimension of the system. If $D$ is sufficiently large, one has
$C(t)\simeq\langle\psi|e^{iHt}{\cal O}e^{-iHt}{\cal O}|\psi\rangle.$ (S4)
The accuracy of Eq. (S4) can be further improved by taking an average over $N_{p}$ different realizations of Haar random states $|\psi\rangle$, and
$\varepsilon(t)\sim\frac{1}{\sqrt{DN_{p}}}.$ (S5)
Employing an auxiliary state
$|\psi_{{\cal O}}\rangle={\cal O}|\psi\rangle,$ (S6)
Eq. (S4) can be written as
$C(t)=\langle\psi(t)|{\cal O}|\psi_{{\cal O}}(t)\rangle,$ (S7)
where
$|\psi(t)\rangle=e^{-iHt}|\psi\rangle,\quad|\psi_{{\cal O}}(t)\rangle=e^{-iHt}|\psi_{{\cal O}}\rangle.$ (S8)
The time evolution of the two states can be calculated by iterating in real time, and we use Chebyshev polynomial algorithm in this paper.
Figure S1: DQT results for the autocorrelation function of the fast-mode observable $\mathcal{O}_{q=\pi}$ in the tilted field Ising model (TFIM) with $B_{z}=0.5$. Left: Results for different system sizes $L$ averaged over $N_{p}=2^{28-L}$ different realizations of Haar-random states. Right: Results for fixed system size $L=18$ and different $N_{p}$. The solid line indicates the exact diagonalization (ED) result as comparison.
Further examples in the tilted-field Ising model
@@
$\mathcal{B}$ $\propto\sum_{k}J\sigma_{k}^{x}\sigma_{k+1}^{x}-\sigma_{k}^{z},$ (S9)
$\mathcal{J}_{E}$ $=B_{x}\sum_{k}\sigma_{k}^{y}\left(\sigma_{k+1}^{z}-\sigma_{k-1}^{z}\right),$ (S10)
with $J=1.05$, for which we compute $n_{\max}=40$ and $n_{\max}=44$ Lanczos coefficients respectively. In Fig. S2 we depict the LP curves for both observables for both $\psi_{0}(t)$ and $\psi_{1}(t)$ as well as the area estimator $\mathcal{A}_{R}$. For the two-local observable $\mathcal{B}$ we find a suitably converged area at $R=18$ and a necessary order $D=3$. For the energy current we have $(R,D)=(35,1)$. Again, the LP scheme gives rise to approximations that are distinguishable from the DQT data to the naked eye. Here, the functions $\psi_{0}(t)$ and $\psi_{1}(t)$ both only consist of at most 21 and 36 exponentially damped constituents for the two observables respectively.
Figure S2: LP method in the tilted field Ising model (TFIM). The two-local observable $\mathcal{B}$ (left) and energy current $\mathcal{J}_{E}$ (right) in the tilted-field Ising model, given by Eq. (S9) and (S10) respectively. From top to bottom: Autocorrelation function obtained by the LP method and DQT (inset: log-plot), the correlation between the first two Krylov states of the respective observable, again by means of the LP method and DQT (inset: log-plot), the area estimate $\mathcal{A}_{r}$ with convergence position $R$ and the respective Lanczos coefficients $b_{n}$.
Lanczos-Pascal for transport
As an application for the Lanczos-Pascal method we compute an approximate diffusion constant for the energy transport in the tilted-field Ising model with $B_{z}=0.5$ by investigating the low wave number limit of the respective density-wave operator (cf. Eq. (10)). Concretely, we proceed as follows: we determine up to 40 Lanczos coefficients for wave numbers $q_{j}=0.01,\dots,0.5$. We then apply the LP scheme for each $q_{j}$ and fit the resulting curve with an ansatz $C_{q_{j}}(t)=\exp{(-\Gamma_{j}t)}$, see Fig. S3. We then compare the decay constants $\Gamma_{j}$ with the respective wave number $q_{j}$. Performing a further quadratic fit of the former for small $q$ (here: $q=0.01,\dots,0.1$), i.e. $\Gamma=\mathcal{D}q^{2}$, we infer a diffusion constant $\mathcal{D}=1.666$, see Fig. S4. Other analyses in the literature report a diffusion constant in this model $D\approx 1.675$, see [28, 34].
Figure S3: Exponential fits of the LP curves for the energy density-wave operator with wave number $q=0.01$ ($a$) and $q=0.1$ ($b$). The LP parameters are $(R,D)=(25,1)$ and $(23,1)$ respectively.
Figure S4: Transport analysis of the LP method. Depicted are the decay constants $\Gamma_{j}$ from the exponential fits to the LP curves for wave numbers $q=0.01,\dots,0.5$. A quadratic fit of form $\Gamma=\mathcal{D}q^{2}$ for small $q=0.01,\dots,0.1$ finds a diffusion constant $\mathcal{D}=1.666.$